A resistor initially at 300k has a het capacity of 10Jk^-1 it is thermally insulted from its surrounding and is connected to a battery from which it takes 1200J through an electric current calculate the change in entropy of the resistor and the surrounding.


1
Expert's answer
2021-04-06T18:23:50-0400

Solution.

"T_1=300K;"

"C=10JK^{-1};"

"\Delta Q=1200J;"

"\Delta S=\dfrac{\Delta Q}{T};"


"C=\dfrac{\Delta Q}{\Delta T}\implies \Delta T=\dfrac{\Delta Q}{C};"

"\Delta T=\dfrac{1200J}{10JK^{-1}}=120K;"

"T_2=T_1+\Delta T;"

"T_2=300K+120K=420K;"

"\Delta S=\dfrac{1200J}{420K}=2.86JK^{-1};"


Answer:"\Delta S=2.86JK^{-1}," entropy of universe remains constant.